GALOIS MODULE STRUCTURE OF MILNOR K-THEORY IN CHARACTERISTIC P

GANESH BHANDARI NICOLE LEMIRE JAN MINAC JOHN SWALLOW

Abstract. Let E be a cyclic extension of degree p^{n} of a field F of characteristic p. We determine k_{m}E, the Milnor K-groups modulo p, as 𝔽_{p}[Gal(E/F)]-modules for all m ∈ ℕ.